# UVIP: Model-Free Approach to Evaluate Reinforcement Learning Algorithms

### v4

@@ 4 Approximate UVIP
In order to implement the approach described in the previous section, we need to construct empirical estimates for the outer expectation and the one-step transition operator ${\rm P}^{a}$ in (3). While in tabular case this boils down to a straightforward Monte Carlo, in case of infinitely many states we need an additional approximation step. Algorithm 2 contains the pseudocode of Approximate UVIP algorithm. Our main assumption is that sampling from ${\rm P}^{a}(\cdot|x)$ is available for any $a\in\mathsf{A}$ and $x\in\mathsf{X}$. For simplicity we assume that the value function $V^{\pi}$ is known, but it can be replaced by its (lower biased) estimate both in Algorithm 2 and subsequent theoretical results. The proposed algorithm proceeds as follows.
@@
Generate $r^{a}(X_{i}),Y^{\mathsf{x}_{i},a}_{j}\sim{\rm P}^{a}(\cdot|\mathsf{x}_{i})$ for all $i\in[N],$ $j\in[M_{1}+M_{2}],$ $a\in\mathsf{A}$;
$k=1$; while  $\|\widetilde{V}_{k}^{\mathrm{up}}-\widetilde{V}_{k-1}^{\mathrm{up}}\|_{\mathsf{X}_{N}}>\varepsilon$ do
@@
$\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{i})=M_{2}^{-1}\sum_{j=M_{1}+1}^{M_{1}+M_{2}}\max_{a}\Bigl{\{}r^{a}(\mathsf{x}_{i})+\gamma\Bigl{(}\widehat{V}_{k}^{\rm{up}}(Y_{j}^{\mathsf{x}_{i},a})-V^{\pi}(Y_{j}^{\mathsf{x}_{i},a})+M_{1}^{-1}\sum_{l=1}^{M_{1}}V^{\pi}(Y_{l}^{\mathsf{x}_{i},a})\Bigr{)}\Bigr{\}}\,,$
where $\mathsf{X}_{N}=\{\mathsf{x}_{1},\ldots,\mathsf{x}_{N}\}$ are design points, either deterministic or sampled from some distribution on $\mathsf{X}.$ Then the next iterate $\widehat{V}^{\rm{up}}_{k+1}$ is obtained via an interpolation scheme based on the points $\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{1}),\ldots,\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{N})$ such that $\widehat{V}_{k+1}^{\rm{up}}(\mathsf{x}_{i})=\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{i}),$ $i=1,\ldots,N.$ Note that interpolation is needed, since $\widehat{V}_{k+1}^{\rm{up}}$ has to be calculated at the (random) points $Y_{j}^{\mathsf{x}_{i},a}$, which may not belong to the set $\mathsf{X}_{N}$. In tabular case when $|\mathsf{X}|<\infty$ is not large one can omit the interpolation and take $\mathsf{X}_{N}=\mathsf{X}$.
In a more general setting when $(\mathsf{X},\rho_{\mathsf{X}})$ is an arbitrary compact metric space we suggest using an appropriate interpolation procedure. The one described below is particularly useful for our situation where the function to be interpolated is only Lipschitz continuous (due to the presence of the maximum). The optimal central interpolant for a function $f\in\mathrm{Lip}_{\rho_{\mathsf{X}}}(L)$ is defined as
@@
Note that $H_{f}^{\mathrm{low}}(x)\leq f(x)\leq H_{f}^{\mathrm{up}}(x),$ $H_{f}^{\mathrm{low}},H_{f}^{\mathrm{up}}\in\mathrm{Lip}_{\rho_{\mathsf{X}}}(L)$ and hence $I[f]\in\mathrm{Lip}_{\rho_{\mathsf{X}}}(L).$ An efficient algorithm is proposed in Beliakov (2006) to compute the values of the interpolant $I[f]$ without knowing $L$ in advance. The so-constructed interpolant achieves the bound
@@ 5 Theoretical results
Furthermore, if $\widehat{V}_{0}^{\rm{up}}(x)\geq V^{\star}(x)$ for $x\in\mathsf{X}_{N},$ then $\overline{V}_{k}^{\rm{up}}(x)\geq V^{\star}(x)$ for any $x\in\mathsf{X}_{N}$ and $k\in\mathbb{N}$. Hence $\widehat{V}_{k}^{\rm{up}}$ is an upper-biased estimate of $V^{\star}$ for any $k\geq 0$.
@@
A 1.
We suppose that $(\mathsf{X},\rho_{\mathsf{X}})$ and $(\mathsf{A},\rho_{\mathsf{A}})$ are compact metric spaces. Moreover, $\mathsf{X}\times\mathsf{A}$ is equipped with some metric $\rho$, such that $\rho\bigl{(}(x,a),(x^{\prime},a)\bigr{)}=\rho_{\mathsf{X}}(x,x^{\prime})$ for any $x,x^{\prime}\in\mathsf{X}$ and $a\in\mathsf{A}$.
@@
A2 is a reparametrization assumption which is popular in RL, see e.g. Ciosek & Whiteson (2020), Heess et al. (2015), Liu et al. (2018) and the related discussions. This assumption is rather mild, since a large class of controlled Markov chains can be represented in the form of random iterative functions, see Douc et al. (2018).
A 3.
@@
Remark 5.1.
If $|\mathsf{X}|<\infty$ and $|\mathsf{A}|<\infty$, the assumption A4 holds with $\rho_{\mathsf{X}}(x,x^{\prime})=\mathbbm{1}_{\{x\neq x^{\prime}\}}$, $\rho((x,a),(x^{\prime},a^{\prime}))=\mathbbm{1}_{\{(x,a)\neq(x^{\prime},a^{\prime})\}}$ and constants $L_{\psi}=1,\,L_{\max}=R_{\max},L_{\pi}=R_{\max}/(1-\gamma)$.
The condition $L_{\psi}\leq 1$ implies a non-explosive behaviour of the Markov chain $(X_{i})_{i\geq 0}$. This assumption is common in theoretical RL studies, see e.g. Pires & Szepesvári (2016). If $L_{\psi}<1$, the corresponding Markov kernel contracts and there exists a unique invariant probability measure, see e.g. Jarner & Tweedie (2001).
Suppose that for each $k\in[K]$ we use an i.i.d. sample $\boldsymbol{\xi}_{k}=(\xi_{k,1}\ldots,\xi_{k,M_{1}+M_{2}})\sim\mathrm{P}_{\xi}^{\otimes(M_{1}+M_{2})}$ to generate $Y^{x,a}_{j}=\psi(x,a,\xi_{k,j}),j\in[M_{1}+M_{2}]$ and these samples are independent for different $k$. For $\varepsilon>0$, we denote by $\mathcal{N}(\mathsf{X}\times\mathsf{A},\rho,\varepsilon)$ the covering number of the set $\mathsf{X}\times\mathsf{A}$ w.r.t. metric $\rho$, that is, the smallest cardinality of an $\varepsilon$-net of $\mathsf{X}\times\mathsf{A}$ w.r.t. $\rho$. Then $\log\mathcal{N}(\mathsf{X}\times\mathsf{A},\rho,\varepsilon)$ is the metric entropy of $\mathsf{X}\times\mathsf{A}$ and
@@
is the Dudley’s integral. Here $\mathsf{D}\doteq\operatorname{diam}(\mathsf{X}\times\mathsf{A})\doteq\max_{(x,a),(x^{\prime},a^{\prime})\in\mathsf{X}\times\mathsf{A}}\rho((x,a),(x^{\prime},a^{\prime}))$. Recall that $\rho(\mathsf{X}_{N},\mathsf{X})$ defined in (8) is the covering radius of the set $\mathsf{X}_{N}$ w.r.t. $\mathsf{X}$. We now state one of our main theoretical results.
Theorem 5.1.
@@
In the above bound $\lesssim$ stands for inequality up to a constant depending on $\gamma,L_{\max},L_{\psi},L_{\pi},L_{0}$ and $R_{\max}$. A precise dependence on the aforementioned constants can be found in (21) in Appendix.
Proof.
The proof is given in Section A.1. ∎
@@
The precise expression for the constants can be found in (22) in Appendix.
Proof.
The proof is given in Section A.2. ∎
Corollary 5.2.
Let $\mathsf{X}\subseteq[0,1]^{d_{\mathsf{X}}}$, $|\mathsf{A}|<\infty$, and consider $\rho_{\mathsf{X}}(x,x^{\prime})=\|x-x^{\prime}\|$, $\rho\bigl{(}(x,a),(x^{\prime},a^{\prime})\bigr{)}=\|x-x^{\prime}\|+\mathbbm{1}_{\{a\neq a^{\prime}\}}$. Assume that A2 – A4 hold and let $\mathsf{X}_{N}=\{\mathsf{x}_{1},\ldots,\mathsf{x}_{N}\}$ be a set of $N$ points independently and uniformly distributed over $\mathsf{X}$. If additionally $\mathrm{Lip}_{\rho_{\mathsf{X}}}(\widehat{V}^{\rm{up}}_{0})\leq L_{0}$ for some $L_{0}>0$, then for any $k\in\mathbb{N}$ and $\delta\in(0,1/2)$ it holds with probability at least $1-\delta$ that
@@
Precise expression for the constants can be found in (23) in Appendix.
Proof.
The proof is given in Section A.2. ∎
@@
where the constant $\operatorname{C}$ depends on $C_{\mathsf{X},\mathsf{A}},\gamma,L_{\max},L_{\psi},L_{\pi},L_{0}$ and $R_{\max}$. A precise expression for $\operatorname{C}$ can be found in (30) in appendix.
Proof.
The proof is given in Section A.3. ∎
Corollary 5.3.
Recall that $\widehat{V}_{k}^{\rm{up}}$ is an upper biased estimate of $V^{\star}$ in a sense that $\overline{V}_{k}^{\rm{up}}(x)\geq V^{\star}(x)$ provided $\widehat{V}_{0}^{\rm{up}}(x)\geq V^{\star}(x)$ for $x\in\mathsf{X}_{N}$. Together with Theorem 5.2, it implies that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$V^{\pi}(x)\leq V^{\star}(x)\leq\widehat{V}_{k}^{{\rm{up}}}(x)\\ +\sigma_{k}\sqrt{\operatorname{C}\log(\mathrm{e}\vee\sigma_{k}^{-1})\delta^{-1}M_{2}^{-1}}+L_{V}\rho(\mathsf{X}_{N},\mathsf{X})\mathbbm{1}_{\{x\not\in\mathsf{X}_{N}\}},\quad x\in\mathsf{X},$ (13)
where the constant $L_{V}$ is given by (18) in appendix.
Note that bounds of type (13) are known in the literature only in the case of specific policies $\pi$. For example, Wainwright (2019) proves bounds of this type for greedy policies in tabular $\mathsf{Q}$-learning. At the same time, (13) holds for arbitrary policy $\pi$ and general state space.
@@
Proof.
The proof is given in Section A.4. ∎
@@ 6 Numerical results
In this section we demonstrate the performance of Algorithm 2 on several tabular and continuous state space RL problems. Recall that the closer policy $\pi$ is to the optimal one $\pi^{\star}$, the smaller is the difference between $V^{\pi}(x)$ and $V^{\rm{up},\pi}(x)$.
@@
We consider $3$ popular tabular environments: Garnet (Archibald et al. (1995)), Chain (Rowland et al. (2020)) and NRoom (Domingues et al. (2021)). Detailed descriptions of these environments are provided in Appendix B. For each environment we perform $K$ updates of the Value iteration (see Appendix B for details) with known transition kernel ${\rm P}^{a}$. We denote the $k$-th step estimate of the action-value function as $\hat{Q}_{k}(x,a)$ and denote by $\pi_{k}$ the greedy policy w.r.t. $\hat{Q}_{k}(x,a)$. Then we evaluate the policies $\pi_{k}$ with the Algorithm 2 for certain iteration numbers $k$. We omit the approximation step because the state space is small. Experimental details are provided in Table 1 in the appendix. Figure 1 displays the gap between $V^{\pi_{k}}(x)$ and $V^{\rm{up},\pi_{k}}(x)$, which converges to zero as $\pi^{k}$ approaches the optimal policy $\pi^{\star}$.
In the NRoom environment, we first learn a sub-optimal policy $\pi$ using the Value Iteration (VI) algorithm. In the third room, we then replace this policy with a uniformly random policy $\pi_{c}$ with probability $1/2$. As expected, this modification results in a less efficient policy within that specific room, which, in turn, should increase the upper bounds of our estimation. To demonstrate this effect, we compute precise upper bounds using the UVIP algorithm. As shown in Figure 1(bottom), UVIP effectively captures the sub-optimality of the policy in the third room, while displaying only slight changes in value estimates for the other rooms.
Figure 1: The difference between $V^{\rm{up},\pi_{i}}(x)$ and $V^{\pi_{i}}(x)$. X-axis represents states in a discrete environment for all pictures. Each group of three pictures of the same color demonstrates the process of learning the policy from the first iteration to the last. First row: Evaluation of the policies during the process of Value iteration for Garnet (left) and Chain environments (right). The policies are the greedy ones corresponding to $Q_{i}(x,a)$ function at the i-th step. Second row: Comparison of the gap between $V^{\pi}$ and $V^{\rm{up},\pi}$ for the learned policy $\pi$ and the corrupted policy $\pi_{c}$ in the NRoom environment. The color in this plot represents the value of $V^{\rm{up},\pi}-V^{\pi}$.
@@
In all subsequent experiments, we obtain sample points $(\mathsf{x}_{1},\dots,\mathsf{x}_{N})$ in Algorithm 2 from trajectories of the evaluation policy. These points are sufficiently representative (see Kveton & Theocharous (2012),Barreto et al. (2016)) and explore key areas of the state space. We consider the AI Gym CartPole and Acrobot environments (see Brockman et al. (2016)), with their descriptions provided in Appendix B. For CartPole we evaluate the A2C algorithm policy $\pi_{1}$ (Mnih et al. (2016)), linear deterministic policy (LD) $\pi_{2}$ described in Appendix B, and random uniform policy $\pi_{3}$. Figure 2 (left) indicates superior quality of $\pi_{2}$, sort of instability introduced by A2C training in $\pi_{1}$ and low quality of $\pi_{3}$. We also evaluated a policy for Acrobot given by A2C as well as a policy from Dueling DQN (Wang et al. (2016)) (Fig. 2 (right)). From the plots we can conclude that the both policies are good, but far from optimal.
@@
Figure 2: Upper and lower bounds of the three different policies. Left: For CartPole $\pi_{1}$, $\pi_{2}$, $\pi_{3}$ policies, respectively. For horizontal axis we sample single trajectory according to the policy. Right: For Acrobot Dueling DQN and A2C policies, respectively. We evaluated the bounds for 50 first states of the trajectory for each algorithm.
@@
Additionally, we compare policies in TwinRooms environment from rlberry (Domingues et al. (2021)). We obtain two policies $\pi_{1}$ and $\pi_{2}$ after running Kernel- UCBVI(Domingues et al. (2022)) algorithm on $2500$ and $5000$ iteration steps. The result on Figure 3 shows that the policy $\pi_{2}$ have tighter upper bounds after more learning steps, and thus, we can conclude that it has better performance. Also, our upper bounds highlight the regions of the state space which are less studied with our policy.
@@
Figure 3: We illustrate the gap between $V^{up,\pi}$ and $V^{\pi}$ in TwinRooms environment. Color on this plot represents the value of $V^{\rm{up},\pi}-V^{\pi}$. On the left and right hand sides we show this quantity for $\pi_{1}$ and $\pi_{2}$ corresponding. We obtain $\pi_{1}$ and $\pi_{2}$ after $2500$ and $5000$ learning steps of the Kernel-UCBVI algorithm.

### v5

@@ 5 Approximate UVIP
In order to implement the approach described in the previous section, we need to construct empirical estimates for the outer expectation and the one-step transition operator ${\rm P}^{a}$ in (3). While in the tabular case this boils down to a straightforward Monte Carlo, in the case of infinitely many states we need an additional approximation step. Algorithm 2 contains the pseudocode of Approximate UVIP algorithm. Our main assumption is that sampling from ${\rm P}^{a}(\cdot|x)$ is available for any $a\in\mathsf{A}$ and $x\in\mathsf{X}$. For simplicity, we assume that the value function $V^{\pi}$ is known, but it can be replaced by its (lower-biased) estimate both in Algorithm 2 and in subsequent theoretical results. We set $G$ as
$G_{k}(x,a,y)=r^{a}(x)+\gamma\bigl(\widehat{V}_{k}^{\rm{up}}(y)-V^{\pi}(y)+M_{1}^{-1}\sum_{\ell=1}^{M_{1}}V^{\pi}(Y_{\ell}^{x,a})\bigr).$
The proposed algorithm proceeds as follows.
@@
Generate $r^{a}(\mathsf{x}_{i}),Y^{\mathsf{x}_{i},a}_{j}\sim{\rm P}^{a}(\cdot|\mathsf{x}_{i})$ for all $i\in[N]$, $j\in[M_{1}+M_{2}]$, $a\in\mathsf{A}$;
$k=1$; while $\sup\limits_{x\in\mathsf{X}_{N}}|\widetilde{V}_{k}^{\mathrm{up}}(x)-\widetilde{V}_{k-1}^{\mathrm{up}}(x)|>\varepsilon$ do
@@
$\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{i})=M_{2}^{-1}\sum_{j=M_{1}+1}^{M_{1}+M_{2}}\max_{a}\Bigl\{G_{k}(\mathsf{x}_{i},a,Y_{j}^{\mathsf{x}_{i},a})\Bigr\}\,$,
where $\mathsf{X}_{N}=\{\mathsf{x}_{1},\ldots,\mathsf{x}_{N}\}$ are design points, either deterministic or sampled from some distribution on $\mathsf{X}.$ Then the next iterate $\widehat{V}^{\rm{up}}_{k+1}$ is obtained via an interpolation scheme based on the points $\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{1}),\ldots,\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{N})$ such that $\widehat{V}_{k+1}^{\rm{up}}(\mathsf{x}_{i})=\widetilde{V}_{k+1}^{\rm{up}}(\mathsf{x}_{i}),$ $i=1,\ldots,N.$ Note that interpolation is needed since $\widehat{V}_{k+1}^{\rm{up}}$ has to be calculated at the (random) points $Y_{j}^{\mathsf{x}_{i},a}$, which may not belong to the set $\mathsf{X}_{N}$. In the tabular case when $|\mathsf{X}|<\infty$ is not large, one can omit the interpolation and take $\mathsf{X}_{N}=\mathsf{X}$.
In a more general setting, when $(\mathsf{X},\rho_{\mathsf{X}})$ is an arbitrary compact metric space, we suggest using an appropriate interpolation procedure. The one described below is particularly useful for our situation, where the function to be interpolated is only Lipschitz continuous (due to the presence of the maximum). The optimal central interpolant for a function $f\in\mathrm{Lip}_{\rho_{\mathsf{X}}}(L)$ is defined as
@@
Note that $H_{f}^{\mathrm{low}}(x)\leq f(x)\leq H_{f}^{\mathrm{up}}(x),$ $H_{f}^{\mathrm{low}},H_{f}^{\mathrm{up}}\in\mathrm{Lip}_{\rho_{\mathsf{X}}}(L)$, and hence $I[f]\in\mathrm{Lip}_{\rho_{\mathsf{X}}}(L).$ An efficient algorithm is proposed in Beliakov (2006) to compute the values of the interpolant $I[f]$ without knowing $L$ in advance. The so-constructed interpolant achieves the bound
@@ 6 Theoretical Results
Furthermore, if $\widehat{V}_{0}^{\rm{up}}(x)\geq V^{\star}(x)$ for $x\in\mathsf{X}_{N},$ then $\overline{V}_{k}^{\rm{up}}(x)\geq V^{\star}(x)$ for any $x\in\mathsf{X}_{N}$ and $k\in\mathbb{N}$. Hence, $\widehat{V}_{k}^{\rm{up}}$ is an upper-biased estimate of $V^{\star}$ for any $k\geq 0$.
@@
A 1
We suppose that $(\mathsf{X},\rho_{\mathsf{X}})$ and $(\mathsf{A},\rho_{\mathsf{A}})$ are compact metric spaces. Moreover, $\mathsf{X}\times\mathsf{A}$ is equipped with some metric $\rho$ such that $\rho\bigl((x,a),(x^{\prime},a)\bigr)=\rho_{\mathsf{X}}(x,x^{\prime})$ for any $x,x^{\prime}\in\mathsf{X}$ and $a\in\mathsf{A}$.
@@
A2 is a reparametrization assumption which is popular in RL, see e.g. Ciosek and Whiteson (2020), Heess et al. (2015), Liu et al. (2017) and the related discussions. This assumption is rather mild, since a large class of controlled Markov chains can be represented in the form of random iterative functions, see Douc et al. (2018).
A 3
@@
Remark 6.1
If $|\mathsf{X}|<\infty$ and $|\mathsf{A}|<\infty$, the assumption A4 holds with $\rho_{\mathsf{X}}(x,x^{\prime})=\mathbb{I}_{\{x\neq x^{\prime}\}}$, $\rho((x,a),(x^{\prime},a^{\prime}))=\mathbb{I}_{\{(x,a)\neq(x^{\prime},a^{\prime})\}}$ and constants $L_{\psi}=1,\,L_{\max}=R_{\max}$, and $L_{\pi}=R_{\max}/(1-\gamma)$.
The condition $L_{\psi}\leq 1$ implies a non-explosive behavior of the Markov chain $(X_{i})_{i\geq 0}$. This assumption is common in theoretical RL studies, see e.g. Ávila Pires and Szepesvári (2016). If $L_{\psi}<1$, the corresponding Markov kernel contracts and there exists a unique invariant probability measure, see e.g. Jarner and Tweedie (2001).
Suppose that we use an i.i.d. sample $\boldsymbol{\xi}_{k}=(\xi_{k,1}\ldots,\xi_{k,M_{1}+M_{2}})\sim\mathrm{P}_{\xi}^{\otimes(M_{1}+M_{2})}$ for each $k\in[K]$ to generate $Y^{x,a}_{j}=\psi(x,a,\xi_{k,j}),j\in[M_{1}+M_{2}]$ and these samples are independent for different $k$. For $\varepsilon>0$, we denote by $\mathcal{N}(\mathsf{X}\times\mathsf{A},\rho,\varepsilon)$ the covering number of the set $\mathsf{X}\times\mathsf{A}$ w.r.t. metric $\rho$, that is, the smallest cardinality of an $\varepsilon$-net of $\mathsf{X}\times\mathsf{A}$ w.r.t. $\rho$. Then $\log\mathcal{N}(\mathsf{X}\times\mathsf{A},\rho,\varepsilon)$ is the metric entropy of $\mathsf{X}\times\mathsf{A}$, and
@@
is the Dudley’s integral. Here $\mathsf{D}\doteq\max\limits_{(x,a),(x^{\prime},a^{\prime})\in\mathsf{X}\times\mathsf{A}}\rho((x,a),(x^{\prime},a^{\prime}))$. Recall that $\rho(\mathsf{X}_{N},\mathsf{X})$ defined in (8) is the covering radius of the set $\mathsf{X}_{N}$ w.r.t. $\mathsf{X}$. We now state one of our main theoretical results.
Theorem 6.1
@@
In the above bound $\lesssim$ stands for inequality up to a constant depending on $\gamma,L_{\max},L_{\psi},L_{\pi},L_{0}$ and $R_{\max}$. A precise dependence on the aforementioned constants can be found in (21) in the Appendix.
Proof
The proof is given in Section A.1.
@@
The precise expression for the constants can be found in (22) in the Appendix.
Proof
The proof is given in Section A.2.
Corollary 6.2
Let $\mathsf{X}\subseteq[0,1]^{d_{\mathsf{X}}}$, $|\mathsf{A}|<\infty$, and consider $\rho_{\mathsf{X}}(x,x^{\prime})=\|x-x^{\prime}\|$, $\rho\bigl((x,a),(x^{\prime},a^{\prime})\bigr)=\|x-x^{\prime}\|+\mathbb{I}_{\{a\neq a^{\prime}\}}$. Assume that A2 – A4 hold and let $\mathsf{X}_{N}=\{\mathsf{x}_{1},\ldots,\mathsf{x}_{N}\}$ be a set of $N$ points independently and uniformly distributed over $\mathsf{X}$. If additionally $\mathrm{Lip}_{\rho_{\mathsf{X}}}(\widehat{V}^{\rm{up}}_{0})\leq L_{0}$ for some $L_{0}>0$, then for any $k\in\mathbb{N}$ and $\delta\in(0,1/2)$ it holds with probability at least $1-\delta$ that
@@
The precise expression for the constants can be found in (23) in the Appendix.
Proof
The proof is given in Section A.2.
@@
where the constant $\operatorname{C}$ depends on $C_{\mathsf{X},\mathsf{A}},\gamma,L_{\max},L_{\psi},L_{\pi},L_{0}$ and $R_{\max}$. A precise expression for $\operatorname{C}$ can be found in (30) in the Appendix.
Proof
The proof is given in Section A.3.
Corollary 6.3
Recall that $\widehat{V}_{k}^{\rm{up}}$ is an upper-biased estimate of $V^{\star}$ in the sense that $\overline{V}_{k}^{\rm{up}}(x)\geq V^{\star}(x)$ provided $\widehat{V}_{0}^{\rm{up}}(x)\geq V^{\star}(x)$ for $x\in\mathsf{X}_{N}$. Together with Theorem 6.2, it implies that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$V^{\pi}(x)\leq V^{\star}(x)\leq\widehat{V}_{k}^{{\rm{up}}}(x)\\ +\sigma_{k}\sqrt{\operatorname{C}\log(\mathrm{e}\vee\sigma_{k}^{-1})\delta^{-1}M_{2}^{-1}}+L_{V}\rho(\mathsf{X}_{N},\mathsf{X})\mathbb{I}_{\{x\not\in\mathsf{X}_{N}\}},\quad x\in\mathsf{X},$ (13)
where the constant $L_{V}$ is given by (18) in the Appendix.
Note that bounds of the type (13) are known in the literature only in the case of specific policies $\pi$. For example, Wainwright (2019) proves bounds of this type for greedy policies in tabular $\mathsf{Q}$-learning. At the same time, (13) holds for an arbitrary policy $\pi$ and a general state space.
@@
Proof
The proof is given in Section A.4.
@@ 7 Numerical Results
In this section, we demonstrate the performance of Algorithm 2 on several tabular and continuous state-space RL problems. Recall that the closer the policy $\pi$ is to the optimal one $\pi^{\star}$, the smaller is the difference between $V^{\pi}(x)$ and $V^{\rm{up},\pi}(x)$.
@@
We consider $3$ popular tabular environments: Garnet (Archibald et al. (1995)), Chain (Rowland et al. (2020)) and NRoom (Domingues et al. (2021a)). Detailed descriptions of these environments are provided in Appendix B. For each environment, we perform $K$ updates of the Value iteration (see Appendix B for details) with known transition kernel ${\rm P}^{a}$. We denote the $k$-th step estimate of the action-value function by $\hat{Q}_{k}(x,a)$ and denote by $\pi_{k}$ the greedy policy w.r.t. $\hat{Q}_{k}(x,a)$. Then we evaluate the policies $\pi_{k}$ with Algorithm 2 for certain iteration numbers $k$. We omit the approximation step because the state space is small. Experimental details are provided in Table 2 in the appendix. Figure 1 displays the gap between $V^{\pi_{k}}(x)$ and $V^{\rm{up},\pi_{k}}(x)$, which converges to zero as $\pi^{k}$ approaches the optimal policy $\pi^{\star}$.
In the NRoom environment, we first learn a suboptimal policy $\pi$ using the Value Iteration (VI) algorithm. In the third room, we then replace this policy with a uniformly random policy $\pi_{c}$ with probability $1/2$. As expected, this modification results in a less efficient policy within that specific room, which, in turn, should increase the upper bounds of our estimation. To demonstrate this effect, we compute precise upper bounds using the UVIP algorithm. As shown in Figure 1(bottom), UVIP effectively captures the suboptimality of the policy in the third room, while displaying only slight changes in value estimates for the other rooms.
Figure 1: The difference between $V^{\rm{up},\pi_{i}}(x)$ and $V^{\pi_{i}}(x)$. The x-axis represents states in a discrete environment for all pictures. Each group of three pictures of the same color illustrates the process of learning the policy from the first iteration to the last. First row: Evaluation of the policies during the process of Value Iteration for Garnet (left) and Chain environments (right). The policies are the greedy ones corresponding to the function $Q_{i}(x,a)$ at the $i$-th step. Second row: Comparison of the gap between $V^{\pi}$ and $V^{\rm{up},\pi}$ for the learned policy $\pi$ and the corrupted policy $\pi_{c}$ in the NRoom environment. The color in these plots represents the value of $V^{\rm{up},\pi}-V^{\pi}$.
@@
In all subsequent experiments, we obtain sample points $(\mathsf{x}_{1},\dots,\mathsf{x}_{N})$ in Algorithm 2 from trajectories of the evaluation policy. These points are sufficiently representative (see Kveton and Theocharous (2012), Barreto et al. (2016)) and explore key areas of the state space. We consider the OpenAI Gym CartPole and Acrobot environments (see Brockman et al. (2016)), with their descriptions provided in Appendix B. For CartPole, we evaluate the A2C algorithm policy $\pi_{1}$ (Mnih et al. (2016)), the linear deterministic policy (LD) $\pi_{2}$ described in Appendix B, and a random uniform policy $\pi_{3}$. Figure 2 (left) indicates the superior quality of $\pi_{2}$, a certain instability introduced by A2C training in $\pi_{1}$, and the low quality of $\pi_{3}$. We also evaluate a policy for Acrobot given by A2C, as well as a policy from Dueling DQN (Wang et al. (2016)) (Fig. 2 (right)). From the plots, we can conclude that both policies are good but far from optimal.
@@
Figure 2: Upper and lower bounds for three different policies. Left: For CartPole $\pi_{1}$, $\pi_{2}$, $\pi_{3}$ policies, respectively. For the horizontal axis, we sample a single trajectory according to the policy. Right: For Acrobot Dueling DQN and A2C policies, respectively. We evaluate the bounds for the first 50 states of the trajectory for each algorithm.
@@
Additionally, we compare policies in the TwinRooms environment from the rlberry (Domingues et al. (2021a)) library. We obtain two policies $\pi_{1}$ and $\pi_{2}$ after running the Kernel-UCBVI (Domingues et al. (2021b)) algorithm for $2500$ and $5000$ iteration steps, respectively. The results in Figure 3 show that after $5000$ learning steps the policy $\pi_{2}$ has a tighter gap between the lower bound $V^{\pi}$ and the upper bound $V^{\rm{up},\pi}$ on the optimal value function. Also, our upper bounds highlight the regions of the state space that are less studied with our policy.
@@
Figure 3: We illustrate the gap between $V^{\rm{up},\pi}$ and $V^{\pi}$ in the TwinRooms environment. The color in these plots represents the value of $V^{\rm{up},\pi}-V^{\pi}$. On the left and right, we show this quantity for $\pi_{1}$ and $\pi_{2}$, respectively. We obtain $\pi_{1}$ and $\pi_{2}$ after $2500$ and $5000$ learning steps of the Kernel-UCBVI algorithm.
Running time of the UVIP algorithm
To demonstrate that the construction of upper bounds is computationally efficient, we compare the UVIP algorithm with the value function estimation algorithm, Fitted Q-Evaluation (FQE). Specifically, we focus on the running time of these algorithms, ensuring a common convergence criterion is applied to both. Let $V_{k}$} represent the estimates of the lower or upper bound at the $k$-th step of the FQE or UVIP algorithms. Additionally, we select a set of sample points $\mathsf{X}_{N}$ at which convergence will be measured. To this end, we define the quantity
$E_{k}=\sqrt{\sum\limits_{x\in\mathsf{X}_{N}}\left(1-{V_{k-1}(x)\over V_{k}(x)}\right)^{2}}\,.$
We stop iterations of both FQE and UVIP procedures when $E_{k}\leq 0.01$. After training the policy for $K$ steps, we construct an $\varepsilon$-greedy policy and evaluate it using both algorithms. Results on the TwinRooms environment are summarized in Table 1. While the UVIP algorithm converges for all instances, it sometimes exhibits significant variance. In contrast, the FQE algorithm fails to converge within the fixed budget for certain seeds. Both algorithms show similar performance in this task.
Table 1: Comparison of running time (in seconds) between the UVIP and FQE algorithms on the TwinRooms environment. The evaluation is conducted for different policies after a specified number of training steps of the Kernel-UCBVI algorithm (represented in each row). The reported running times are averaged over 5 seeds. In some instances, the FQE algorithm does not converge within the fixed budget of 400 epochs. Therefore, results are presented separately: one set includes all seeds, while another considers only the converged cases.
policy training steps FQE all seeds FQE converged UVIP
$1250$ $73.60\pm 13.26$ $73.60\pm 13.26$ $49.23\pm 66.29$
$2500$ $120.85\pm 94.54$ $65.56\pm 14.24$ $78.36\pm 77.54$
$5000$ $110.06\pm 85.35$ $87.06\pm 65.36$ $82.26\pm 107.77$